Will you also please show me how to solve the system by addtion
5x-3y=13
4x-3y=11
Thank you

Answer by stanbon(75887)

Will you also please show me how to solve the system by addtion
1st: 5x-3y=13
2nd: 4x-3y=11
Subtract 2nd from 1st to get:
3rd: x=2
Substitute that into 1st to solve for y:
5(2)-3y=13
-3y=3
y=-1
Check these answers in 2nd as follows:
4x-3y=11
4(2)-3(-1)=11
8+3=11
11=11
Cheers,
Stan H.